the local theater charges $8 for adult and $5 for children. at the noon performance, there were twice as many children as adult and the total money collected was $576.how many children attended the show?
Answer by checkley77(12569)

6x+3*2x=576
6x+6x=576
12x=576
x=576/12
x=48 adult tickets were sold.
2*48=96 children tickets were sold.
Proof:
6*48+3*96=576
288+288=576
576=576